# Latent Neural Operator for Solving Forward and Inverse PDE Problems

### v1

@@ 3.3 Physics Cross Attention
Correspondingly, the cross-attention in the decoder has a learnable key matrix that represents sample positions in the latent space. Therefore, the decoder in  (3) is written as
@@
In the special case where the input only contains sample positions without physics quantities, our PhCA module is similar to the transformation of physics-aware token in Transolver [23] (only differing in normalizations). In general cases when the inputs inlcude both positions and physics quantities, the two approaches are different. Specifically, in Transolver, the key $\boldsymbol{X}$ in (4) and query $\boldsymbol{P}$ in (5) contain both positions and physics quantities information, whereas in PhCA, the key matrix $\boldsymbol{X}$ in the encoder and the query matrix $\boldsymbol{P}$ in the decoder only depend on the positions of the input and output respectively. The physics quantities are used only to produce the value matrix. Such decoupling property of values and locations in query and key enables LNO to predict values at different positions from input positions.
@@ 4 Experiments
We first evaluate the accuracy for the forward problems on the six popular public benchmarks and for the inverse problem using Burger’s equation. Then, we measure the efficiency in term of the number of model parameters, GPU memory usage and the training time. Finally, we establish ablation study on the choices of different types of attention mechanisms and sample sizes in latent space. Additional experiments on scaling, sampling rate, and sharing weights between encoder and decoder are included in the appendix.
@@ 4.1 Accuracy for the forward problems
Table 1 shows a comprehensive comparison with previous works who reported their performance on the same benchmarks. Our LNO achieves best accuracy on four benchmarks: Darcy, NS2d, Elasticity and Pipe. On the Airfoil and Plasticity benchmarks, our LNO achieves performance close to the SOTA with only half computational cost discussed in Section 4.3. These results demonstrate the effectiveness of approximating the kernel integral operator in the latent space through PhCA module.
Table 1: Prediction error on the six forward problems. The best result is in bold. We reproduce the Transolver by implementing codes independently (marked with $\ast$) besides the results claimed in the original paper[23]. The column labeled D.C. indicates whether the observation positions and prediction positions are decoupled. Standard deviations are computed based on 5 independent trials.
@@ 4.2 Accuracy for the inverse problem
To demonstrate the flexibility of LNO, we design an inverse problem. Given a set of partially observed solution $u(x)$, the aim is to recovery the complete $u(x)$ in a larger domain. Specifically, we test on 1D Burgers’ equation
@@
The objective of this inverse problem is to reconstruct the complete solution $u(x)$ in the whole spatiotemporal domain $(x,t)\in[0,1]\times[0,1]$ based on sparsely random-sampled or fixed observation in the sub-spatiotemporal domain $(x,t)\in[0,1]\times[0.25,0.75]$, as illustrated in Figure 6 in the appendix.
Instead of naive approach which directly learns the mapping from partially observed samples in the subdomain to the complete solution in the whole domain, we propose a two-stage strategy. First, we train the model as a completer to interpolate the solution in the subdomain. Then, we train the model as a propagator to extrapolate from the subdomain to the whole domain. Since, the observation and the prediction samples locate in different positions, only the models with decoupling property can be used as the completer and propagator.
We compare the performance of our LNO with DeepONet[13] and GNOT[21] in two stages respectively. The results in Table 2 and Table 3 indicate that i) in the first stage, LNO performs significantly better than DeepONet and GNOT at different random observation ratio; ii) in the second stage, LNO performs better when the complete solution in the subdomain approaches the ground truth, and the performance degrades as the reconstruction error of the complete solution in the subdomain increases. This is reasonable since we train the propagator based on the ground truth in the sub domain. When there is a significant discrepancy between the ground truth and the prediction of completer, the model faces challenging distribution shifts. Nevertheless, when LNO is used both as the completer and propagator, it achieves the most accurate results for solving this inverse problem.
@@
Table 3: The reconstruction error of different propagators in the 2nd stage of the inverse problem. Propagators trained to reconstruct the solution in the whole domain based on the ground truth (G.T.) of the subdomain and are tested using the output of different completers. Relative MAE of $t=0$ and $t=1$ is reported.
@@
Table 4: Comparison of efficiency between LNO and Transolver on the six forward problem benchmarks. Three metrics are measured: number of parameters, cost of GPU memory and cost of training time per epoch.
@@ 4.3 Efficiency
We demonstrate improvement on efficiency of our proposed LNO compared to Transolver, which was considered the most efficient Transformer-based PDE solvers[23]. Results in Table 4 indicate that our method has fewer parameters, lower memory consumption, and faster training speed. Roughly speaking, LNO reduces the parameter count by an average of 30% and memory consumption by 50% compared to Transolver across all benchmarks, also with a 1.8x training speed up.
The improvement of LNO stems from the idea of solving PDEs in the latent space. Given $N$ input points in real-world space, if both of Transolver and LNO are consisted of $L$ Transformer blocks, the Transolver, which applies Physics-Attention in each block, has the time complexity of $O(LMN+LM^{2})$, where $M$ is the number of slices. Our LNO, which applies Physics-Cross-Attention only in the encoder and decoder, has the time complexity of $O(MN+LM^{2})$, where $M$ is the sample size in latent space.
@@ 4.4 Ablation Study
Decoupling the values and locations as well as the observation positions and prediction positions not only guarantees the flexibility but also improves the accuracy. To study the effect of decoupling property, we replaces the PhCA in LNO by Physics-Attention proposed in Transolver[23]. In this way, the query and keys are calculated from the same embedding of both ${pos}_{in}$ and ${val}_{in}$ of input sequence. The ${pos}_{out}$ which not presents in the input sequence can not generate queries and keys. So the observation positions and prediction positions must be kept the same. We compare the performance of LNO with PhCA or Physics-Attention on the forward problem benchmarks, and the results in the first and last rows of Table 5 verify the necessity of decoupling.
@@
We study the impact on accuracy using different attention mechanisms in the latent space. We replace the scaled dot-product attention in LNO with Galerkin-type attention used in Galerkin Transformer[20], efficient attention[18] used in GNOT[21], and Nystrom attention[19] used in ONO[29]. The comparison results in Table 5 show that quadratic attention such as scaled dot-product attention[16] is more expressive than linear attention, which helps improve the accuracy. Many previous methods have employed different linear attentions[20, 21, 29] to reduce the computational costs at the expense to reduce the accuracy. Since the length of the latent sample sequence is much smaller than the inputs and outputs in real-world space, therefore, applying quadratic attention in latent space does not incur significant overhead.
Table 5: The ablation results using different attention implementation. P.A. stands for Physics-Attention. E.A. stands for Efficient Attention, N.A. stands for Nystrom Attention and G.A. stands for Galerkin-type Attention.
@@
Sample size in latent space
We investigate the impact of sample size in latent space on LNO for solving both the forward and inverse problems. Results are shown in Figure 3. In the forward problems, as the sample size $M$ in latent space increases, the performance of LNO gradually improves to a saturation level, followed by a gentle decline on the benchmarks of Airfoil, Elasticity, Plasticity, and Pipe. However, on benchmarks of Darcy and NS2d, the performance of LNO continues to improve till $M=512$. This may be attributed to the former set of benchmarks containing only location information, while the latter set involves both location and physics quantity information, which requires more samples to represent. In the inverse problem, the performance of completer and propagator are similar to the trends in the forward problems.
Figure 3: Sample size in latent space v.s. accuracy. (a) Forward problems on Airfoil, Elasticity, Plasticity, Pipe. (b) Forward problems on Darcy, NS2d. (c) Inverse Problem.
@@ 5 Conclusions
The paper explores the feasibility of characterizing PDEs in a learnable latent space. We present Physics-Cross-Attension (PhCA) for encoding inputs into the latent space and for decoding outputs. We build the Latent Neural Operator (LNO) based on PhCA and validate its flexibility, efficiency and accuracy for solving both forward and inverse problems. In this work, no prior is imposed on the latent space, but, in future, it is worthy to explore whether physical priors, e.g. symmetry or disentanglement, can help gain further improvements, and explore generalization ability of PhCA crossing domains and re-usability for multiple types of PDEs.
Appendix A Scaling Experiments
@@ Appendix B Observation Ratio and Locations for the Inverse PDE Task
Figure 8: Accuracy of solving inverse problem at different temporal and spatial sampling intervals. (a) Accuracy of LNO as completer. (b) Accuracy of LNO as propagator based on the results of corresponding completer.
Appendix C Additional Ablation Study
We conduct additional ablation study to explore the performance of LNO when ${W}_{1}\neq{W}_{2}$ in (4) and (5). In this way, we replace the linear projection $W_{1}$ and $W_{2}$ with two individual MLPs. Results of comparison between LNO using shared MLPs and LNO using non-shared MLPs in PhCA encoder and decoder are shown in Table 6. It can be found that LNO performs better on most benchmarks (except Plasticity) under the shared situation.
Table 6: The ablation results on six forward problem benchmarks. Relative L2 is recorded. Smaller values represent better performance. The better result is in bold.
@@ Appendix C Additional Ablation Study
LNO(non-shared) 0.54 10.67 0.53 1.09 0.26 0.36
LNO 0.49 8.45 0.51 0.52 0.29 0.26
@@ Appendix D Full Implementation Details
For the forward problems, we train the models using relative L2 error for 500 epochs. AdamW optimizer[44] and OneCycleLr scheduler[45] are applied with initial learning rate at $10^{-3}$. We reproduced the results of Transolver[23] following implementations in appendix B.3 in [23] and also excerpt the results of other methods from Table 2 in [23]. The data splits are the same as in appendix B.1 in [23].
For the inverse problem, we train the models using mean squared error and StepLr scheduler as well. For DeepONet[13], we use 4 Transformer blocks with 128-dimensional tokens as the branch net and 3 fully-connected layers as the trunk net. For GNOT[21], we set the number of Transformer blocks as 4, dimension of tokens as 96, and number of experts as 1. We use 4096 samples for training, 128 samples for validating and 128 samples for testing.

### v3

@@ 3.3 Physics-Cross-Attention
Correspondingly, the cross-attention in the decoder has a learnable key matrix that represents positions of samples in the latent space. Therefore, the cross-attention of the decoder in  (3) is written as
@@
In the special case where the input only contains sample positions without physics quantities, our PhCA module is similar to the transformation of physics-aware token in Transolver [23] (only differing in normalization). In general cases when the inputs include both positions and physics quantities, the two approaches are different. Specifically, in Transolver, the key $\hat{\boldsymbol{X}}$ in (4) and query $\boldsymbol{P}$ in (5) contain both positions and physics quantities information, whereas in PhCA, the key matrix $\hat{\boldsymbol{X}}$ in the encoder and the query matrix $\boldsymbol{P}$ in the decoder only depend on the positions of the input and output respectively. The physics quantities are used only to produce the value matrix $\boldsymbol{V}$ in attention mechanism. Such decoupling property of values and locations in query and key enables LNO to predict values at different positions from input positions.
@@ 4 Experiments
We first evaluate the accuracy for the forward problems on the six popular public benchmarks and for the inverse problem using Burger’s equation. Then, we measure the efficiency in terms of the number of model parameters, the GPU memory usage and the training time. Next, we assess the generalization of our model concerning the number $N_{in}$ of observation samples and the number $N_{out}$ of prediction positions. Finally, we establish a series of ablation studies including choices of different attention mechanisms, strategy of sharing weights and sampling numbers in the latent space. Additional experiments on scaling, sampling rate, and significance are presented in the appendix.
@@ 4.1 Accuracy for the forward problems
Table 1 shows a comprehensive comparison with previous works that reported their performance on the same benchmarks. Our LNO achieves the best accuracy on four benchmarks: Darcy, NS2d, Elasticity and Pipe. On the Airfoil and Plasticity benchmarks, our LNO achieves performance close to the SOTA with only half computational cost as discussed in Section 4.3. These results demonstrate the effectiveness of approximating the kernel integral operator in the latent space after transformation from the real-world geometric space through the PhCA module.
Table 1: Prediction error on the six forward problems. The best result is in bold. $\ast$" means that the results of the method are reproduced by ourselves. "///" means that the method can not handle this benchmark. The column labeled D.C. indicates whether the observation positions and prediction positions are decoupled.
@@ 4.2 Accuracy for the inverse problem
To demonstrate the flexibility of LNO, we design an inverse problem. Given the partially observed solution $u(x)$, the aim is to recover the complete $u(x)$ in a larger domain. Specifically, we test on 1D Burgers’ equation
@@
The objective of this inverse problem is to reconstruct the complete solution $u(x)$ in the whole spatiotemporal domain $(x,t)\in[0,1]\times[0,1]$ based on sparsely random-sampled or fixed-sampled observation in the sub-spatiotemporal domain $(x,t)\in[0,1]\times[0.25,0.75]$, as illustrated in Figure 6 in the appendix.
Instead of the naive approach which directly learns the mapping from partially observed samples in the subdomain to the complete solution in the whole domain, we propose a two-stage strategy inspired by inpainting[47, 48] and outpainting[49, 50]. First, we train the model as a completer to interpolate the sparsely sampled points in the subdomain $[0,1]\times[0.25,0.75]$ (left sub-figure in Figure 6) to predict all densely and regularly sampled points in the same subdomain (middle sub-figure in Figure 6). Then, we train the model as a propagator to extrapolate the results of the completer from the subdomain to the whole domain $[0,1]\times[0,1]$ (right sub-figure in Figure 6). Since the observation and the prediction samples are located in different positions, only the models with decoupling properties can be used as the completer and propagator.
We compare the performance of our LNO with DeepONet[13] and GNOT[21] in two stages respectively. The results in Table 2 and Table 3 indicate that i) in the first stage, LNO performs significantly better than DeepONet and GNOT at different random observation ratios; ii) in the second stage, LNO performs better when the complete solution in the subdomain approaches the ground truth, and the performance degrades as the reconstruction error of the complete solution in the subdomain increases. This is reasonable since we train the propagator based on the ground truth in the subdomain. When there is a significant discrepancy between the ground truth and the prediction of the completer, the model faces challenging distribution shifts. Nevertheless, when LNO is used both as the completer and propagator, it achieves the most accurate results for solving this inverse problem.
@@
Table 3: The reconstruction error of different propagators in the 2nd stage of the inverse problem. Propagators are trained to reconstruct the solution in the whole domain based on the ground truth (G.T.) of the subdomain and are tested using the output of different completers. Relative MAE of $t=0$ and $t=1$ is reported.
@@
Table 4: Comparison of efficiency between LNO and Transolver on the six forward problem benchmarks. Three metrics are measured: the number of parameters, the cost of GPU memory and the cost of training time per epoch.
@@ 4.3 Efficiency
We demonstrate improvement in the efficiency of our proposed LNO compared to Transolver, which was considered the most efficient Transformer-based PDE solvers[23]. Results in Table 4 indicate that our method has fewer parameters, lower memory consumption, and faster training speed. Roughly speaking, LNO reduces the parameter count by an average of 30% and memory consumption by 50% compared to Transolver across all benchmarks, also with a 1.8x training acceleration.
The improvement of LNO stems from the idea of solving PDEs in the latent space. Given $N$ input points in the real-world geometric space, if both Transolver and LNO consist of $L$ Transformer blocks, the Transolver, which applies Physics-Attention in each block, has the time complexity of $O(LMN+LM^{2})$, where $M$ is the number of slices. Our LNO, which applies Physics-Cross-Attention only in the encoder and decoder, has the time complexity of $O(MN+LM^{2})$, where $M$ is the sampling numbers in latent space.
4.4 Generalization
To validate the generalization of our model concerning the number of discretization points in the input and output, we train the model in a uniform low-resolution scenario and perform zero-shot inference in other higher-resolution scenarios. Specifically, we downsample the Darcy dataset with a resolution of $421\times 421$ to $241\times 241$, $211\times 211$, $141\times 141$, $85\times 85$, $61\times 61$ and $43\times 43$. Models are trained on the $43\times 43$ dataset and tested on the other datasets. The comparison results of our model and other methods shown in Table  5 show that our model exhibits great generalization capability regarding the number of sampling points, consistently outperforming other methods in accuracy on test sets with a different number of discretization points from the training set.
Table 5: Prediction errors on Darcy with different resolutions. Models are trained on $43\times 43$ and tested on higher resolutions. Relative L2 errors are reported. The best results are in bold.
Model $61\times 61$ $85\times 85$ $141\times 141$ $211\times 211$ $241\times 241$
FNO 0.1164 0.1797 0.2679 0.3160 0.3631
ONO 0.0204 0.0259 0.0315 0.0349 0.0386
GNOT 0.0256 0.0275 0.0294 0.0305 0.0317
Transolver 0.0239 0.0240 0.0258 0.0279 0.0290
LNO(Ours) 0.0179 0.0192 0.0214 0.0229 0.0244
@@ 4.5 Ablation Study
Decoupling the values and locations as well as the observation positions and prediction positions not only guarantees flexibility but also improves accuracy. To study the effect of decoupling property, we replace the PhCA in LNO with Physics-Attention proposed in Transolver[23]. In this way, the query and keys are calculated from the same embedding of both ${pos}_{in}$ and ${val}_{in}$ of the input sequence. The ${pos}_{out}$ which is not present in the input sequence can not generate queries and keys. So the observation positions and prediction positions must be kept the same. We compare the performance of LNO with PhCA or Physics-Attention on the forward problem benchmarks, and the results in the first and last rows of Table 6 verify the necessity of decoupling.
@@
We study the impact on accuracy using different attention mechanisms in the latent space. We replace the scaled dot-product attention in LNO with Galerkin-type attention used in Galerkin Transformer[20], efficient attention[18] used in GNOT[21], and Nystrom attention[19] used in ONO[31]. The comparison results in Table 6 show that quadratic attention, i.e., the scaled dot-product attention[16], is more expressive than linear attention, which helps improve the accuracy. Many previous methods have employed different linear attentions[20, 21, 31] to reduce the computational costs at the expense of reducing the accuracy. Since the length of the latent sequence is much smaller than the inputs and outputs in real-world geometric space, applying quadratic attention in latent space does not incur significant overhead.
Sharing parameters
We explore the performance of LNO when ${W}_{1}\neq{W}_{2}$ in (4) and (5). In this way, we replace the linear projection $W_{1}$ and $W_{2}$ with two individual MLPs. Results of the comparison between LNO using shared MLPs and non-shared MLPs in the PhCA encoder and decoder are shown in Table 7. It can be found that LNO performs better on most benchmarks (except Plasticity) under the shared situation.
Sampling numbers in latent space
We investigate the impact of hypothetical sampling position number in the latent space on LNO for solving both the forward and inverse problems. Results are shown in Figure 3. In the forward problems, as the sampling number $M$ in the latent space increases, the performance of LNO gradually improves to a saturation level, followed by a gentle decline on the benchmarks of Airfoil, Elasticity, Plasticity, and Pipe. However, on benchmarks of Darcy and NS2d, the performance of LNO continues to improve till $M=512$. This may be attributed to that the former set of benchmarks containing only location information, while the latter set involves both location and physics quantity information, which requires more sampling positions to represent. In the inverse problem, the performance of the completer and propagator are similar to the trends in the forward problems.
Table 6: The ablation results using different attention implementations. P.A. stands for Physics-Attention. E.A. stands for Efficient Attention, N.A. stands for Nystrom Attention and G.A. stands for Galerkin-type Attention.
@@
Table 7: The ablation results of strategy of sharing weights on six forward problem benchmarks. Relative L2 is recorded. Smaller values represent better performance. The better result is in bold.
Model Relative L2(×$\times 10^{-2}$)
Darcy NS2d Airfoil Elasticity Plasticity Pipe
LNO(non-shared) 0.54 10.67 0.53 1.09 0.26 0.36
LNO 0.49 8.45 0.51 0.52 0.29 0.26
Figure 3: Sampling numbers in latent space v.s. Accuracy. (a) Forward problems on Airfoil, Elasticity, Plasticity and Pipe. (b) Forward problems on Darcy and NS2d. (c) Inverse Problem.
@@ 5 Conclusions
The paper explores the feasibility of characterizing PDEs in a learnable latent space. We present Physics-Cross-Attension (PhCA) for encoding inputs into the latent space and for decoding outputs. We build the Latent Neural Operator (LNO) based on PhCA and validate its flexibility, efficiency and accuracy for solving both forward and inverse problems. We leave the generalization ability and re-usability of PhCA across multiple types of PDEs as another open research topic.
Our work has also some limitations. For time-dependent equations, we decode the states at intermediate time steps into geometric space to compute the loss. Implementing additional training strategies may allow us to conduct all operations of the dynamic evolution process directly in the latent space, which potentially enhances efficiency. Also, in our work, no prior is imposed on the latent space. However, in the future, it is worth exploring whether physical priors, e.g., symmetry, can help gain further precision improvement.
Appendix A Scaling
@@ Appendix B Observation Ratio and Locations for the Inverse PDE Task
Figure 8: Accuracy of solving the inverse problem at different temporal and spatial sampling intervals. (a) Accuracy of LNO as completer. (b) Accuracy of LNO as propagator based on the results of the corresponding completer.
Appendix C Significance
We conduct repeated experiments on both the forward and inverse problems, initializing the weights of our model and the reproduced models with different random seeds each time. We report the mean and standard deviation of these experiments to demonstrate the significance of our experimental results, as shown in Figure 8 and Figure 9.
Table 8: The mean and standard deviations (computed based on 5 independent trials) of different models in solving the forward problem.
@@ Appendix C Significance
F-FNO[45] 0.75($\pm$0.01) 11.51($\pm$0.25) 0.60($\pm$0.03) 1.85($\pm$0.03) 0.27($\pm$0.02) 0.68($\pm$0.02)
U-FNO[46] 1.28($\pm$0.02) 17.15($\pm$0.29) 1.19($\pm$0.05) 2.13($\pm$0.03) 0.41($\pm$0.03) 0.62($\pm$0.02)
GNOT[21] 1.04($\pm$0.02) 13.40($\pm$0.28) 0.75($\pm$0.02) 0.88($\pm$0.03) 3.19($\pm$0.07) 0.45($\pm$0.02)
Transolver[23] 0.58($\pm$0.01) 8.79($\pm$0.17) 0.47($\pm$0.02) 0.62($\pm$0.02) 0.12($\pm$0.02) 0.31($\pm$0.01)
LNO(Ours) 0.49($\pm$0.01) 8.45($\pm$0.22) 0.51($\pm$0.05) 0.52($\pm$0.03) 0.29($\pm$0.03) 0.26($\pm$0.03)
Table 9: The mean and standard deviations (computed based on 3 independent trials) of different models in solving the inverse problem. Relative MAEs are expressed as percentages. The second to the second-to-last columns correspond to the situation where models act as completers under different observation rates, while the last column corresponds to the situation where models act as propagators.
Model 20% 10% 5% 1% 0.5% propagator
DeepONet[13] 2.37($\pm$0.18) 2.52($\pm$0.04) 2.77($\pm$0.10) 3.18($\pm$0.03) 4.70($\pm$0.31) 7.42($\pm$0.37)
GNOT[21] 1.23($\pm$0.07) 1.33($\pm$0.06) 1.63($\pm$0.04) 1.75($\pm$0.02) 2.53($\pm$0.08) 5.43($\pm$0.10)
LNO(Ours) 0.60($\pm$0.03) 0.73($\pm$0.11) 0.78($\pm$0.04) 1.26($\pm$0.09) 1.97($\pm$0.04) 3.46($\pm$0.21)
@@ Appendix D Full Implementation Details
For the forward problems, we train the models using relative L2 error for 500 epochs. AdamW optimizer[51] and OneCycleLr scheduler[52] are applied with initial learning rate at $10^{-3}$. We reproduce the results of Transolver[23] following implementations in appendix B.3 in [23] and also excerpt the results of other methods from Table 2 in [23]. The data splits are the same as in Appendix B.1 in [23].
For the inverse problem, we train the models using mean squared error and StepLr scheduler. For DeepONet[13], we use 4 Transformer blocks with 128-dimensional tokens as the branch net and 3 fully-connected layers as the trunk net. For GNOT[21], we set the number of Transformer blocks as 4, the dimension of tokens as 96, and the number of experts as 1. We use 4096 samples for training, 128 samples for validating and 128 samples for testing.